6. "A researcher analyzed the results of an experiment and found that the obtained t-value (on a t-test of independent means) was 1.29, with a total of 25 children in group 1 and 30 children in group 2. Use a table of critical values and discuss whether the null hypothesis can or cannot be rejected. (Hint: Use Table 8.6 and 30 as the degree of freedom to support your decision.)"

14. For the following set of 10 scores, compute the range, the standard deviation, and the variance. 5, 7, 3, 4, 5, 6, 7, 2, 5, 3.

Range:

Standard deviation (3 digits after decimal point):

Variance (1 digit after decimal point):
Answer #1

14)

arrange the numbers in ascending order

2,3,3,4,5,5,5,6,7,7

Range = (LArge - small)number
= ( 7 -2) =5

variance = sum of((x - mean)^2/n-1)

mean = sum of all number/ toatl number'
=47/10 = 4.7

variance= (5 - 4.7)^2 + (7 - 4.7)^2 + (3 - 4.7)^2 + (4 - 4.7)^2 + (5 - 4.7)^2 + (6 - 4.7)^2 + (7 - 4.7)^2 + (2 - 4.7)^2 + (5 - 4.7)^2 + (5 - 4.7)^2/9
= 2.9

std.deviation = sqrt(2.9)
= 1.703